OpenStudy (anonymous):

Ken has a pole that is 26 inches long. He wishes to cut it into two pieces so that one piece will be 8 inches longer than the other. How long must the shorter piece be?

OpenStudy (nurali):

One piece = x second piece = x + 8 x + x + 8 = 26 2x = 26 - 8 2x = 18 x = 9 Short piece is 9"

OpenStudy (nurali):

Welcome to Openstudy